# PixWizard: Versatile Image-to-Image Visual Assistant with Open-Language Instructions

### v1

@@ Abstract
This paper presents a versatile image-to-image visual assistant, PixWizard, designed for image generation, manipulation, and translation based on free-from language instructions. To this end, we tackle a variety of vision tasks into a unified image-text-to-image generation framework and curate an Omni Pixel-to-Pixel Instruction-Tuning Dataset. By constructing detailed instruction templates in natural language, we comprehensively include a large set of diverse vision tasks such as text-to-image generation, image restoration, image grounding, dense image prediction, image editing, controllable generation, inpainting/outpainting, and more. Furthermore, we adopt Diffusion Transformers (DiT) as our foundation model and extend its capabilities with a flexible any resolution mechanism, enabling the model to dynamically process images based on the aspect ratio of the input, closely aligning with human perceptual processes. The model also incorporates structure-aware and semantic-aware guidance to facilitate effective fusion of information from the input image. Our experiments demonstrate that PixWizard not only shows impressive generative and understanding abilities for images with diverse resolutions but also exhibits promising generalization capabilities with unseen tasks and human instructions. The code and related resources are available at https://github.com/AFeng-x/PixWizard
@@ 1 Introduction
  1. 1.
    Task Unification: Given the diverse nature of vision tasks and data formats, ranging from pixels and coordinates to binary masks and categories, it is challenging to find a unified representation. We observe that most vision tasks can be framed as image-to-image translation problems. For tasks not naturally suited to image output, we first learn to generate their visualizations and then apply post-processing to convert them into the desired formats. This approach marks an initial step toward building a versatile visual assistant.
  2. 2.
    Data Construction: We aim to leverage and integrate the remarkable diversity of tasks and data in the visual domain. To achieve this, we have built a comprehensive training set with a total of 30 million data points, enabling our model to support five main capabilities: (i) Image generation, which includes text-to-image generation, controllable generation, inpainting, and outpainting; (ii) Image editing; (iii) Image restoration, covering tasks such as deraining, desnowing, deblurring, super-resolution, and more; (iv) Image grounding, which involves locating objects based on user prompts; and (v) Dense image prediction, which includes depth estimation, surface normal estimation, pose estimation, semantic segmentation, and image-to-canny/HED/sketch/Line-Art conversions.
  3. 3.
    Architecture Design: For a robust visual assistant, the architecture and scalability of the foundational model are crucial. In this work, we use the flow-based Diffusion Transformer (DiT) (Ma et al., 2024) as our base model. Compared to traditional diffusion models, flow-based models offer greater versatility and stability in modeling data distributions by learning conditional velocity fields. The DiT architecture further enhances scalability and is well-suited for incorporating conditional information. Building on this foundation, we introduce several unique and effective components. Specifically, we extend the dynamic partitioning and padding scheme to enable the model to handle input images of any resolution, closely aligning with human perceptual processes. Additionally, we introduce structure-aware and semantic-aware guidance, allowing the model to effectively follow multimodal instructions (images and open-ended user prompts) for a variety of manipulations.
@@ 2 Omni Pixel-to-Pixel Instruction-Tuning Dataset
To empower our image-to-image visual assistant with the capabilities on comprehensive image generation, manipulation, and translation, we first compile a multi-task, multi-model training dataset for visual instruction tuning, consisting of 30 million instances across seven primary domains. To our knowledge, this dataset is the largest, most diverse, and easiest-to-use image-instruction-image triplet dataset. It is constructed from both open-source and in-house datasets, filtered with the assistance of MLLMs and manual review, as illustrated in Fig. 1.
Image Restoration. We incorporate low-level data to restore images degraded by various environmental or technical factors. This section utilizes a wide range of open-source datasets covering key restoration tasks, including (1) Denoising, (2) Deraining, (3) Demoireing, (4) Dehazing, (5) Deblurring, (6) Desnowing, (7) Deshadowing, (8) Low-Light Enhancement, (9) Flash Reflection Removal, (10) Face Restoration, (11) Watermark Removal, and (12) Super Resolution. Since both the inputs and outputs are inherently defined in the RGB space, these tasks can be seamlessly unified by our PixWizard model without extra any transformations. All open-source datasets we use are provided in Sec. B.1.
@@
Inpainting, in the context of image generation, involves filling in missing parts of an image with new or modified content. To create inpainting instances, we apply random black or white masks to different regions of the original images. These masks come in various shapes, including circles, rectangles, and free-form patterns, and are randomly placed on the images, resulting in a wide range of occluded areas for inpainting.
@@ 3.2 Architecture
Semantic-Aware Guidance. Besides recognizing structural features, semantic information is also crucial. Therefore, we use CLIP L/14-336 (Radford et al., 2021) to obtain semantic image embeddings. We introduce two zero-initialized attention mechanism within the PixWizard block, which allows the latent target image tokens to query information from the keys and values of the conditions. Specifically, we employ a zero-initialized gating mechanism to gradually inject conditional image and text information into the token sequences. Given target image queries $Q_{i}$, keys $K_{i}$, and values $V_{i}$, along with text instruction keys $K_{t}$ and values $V_{t}$, and conditional image keys $K_{ci}$ and values $V_{ci}$, the final attention output is formulated as:
@@ 3.3 Any Resolution
PixWizard inherits the dynamic partitioning and padding scheme introduced by (Zhuo et al., 2024), enabling the model to handle images of any resolution and aspect ratio during both fine-tuning and inference. However, in practice, the required resolutions for different tasks can vary significantly. To support more flexible handling of arbitrary resolutions and to preserve the original resolution of each image as much as possible, we use $[512^{2},768^{2},1024^{2}]$ as resolution centers to generate a set of candidate patch partitions. During training, we group data items with similar resolutions into the same bucket to ensure that the longest and shortest sequences within each batch are not too different in length, thereby minimizing the use of padding tokens and improving training efficiency. During inference, by incorporating NTK-Aware Scaled RoPE (Peng & Quesnelle, 2023) and sandwich normalization, PixWizard also demonstrates exceptional resolution extrapolation.
@@ 4.1 First Section Results
Dense image prediction tasks are evaluated on the three vision tasks: ADE20k (2017b) for semantic segmentation, NYUv2 (2012) and SUNRGB-D (2014) for monocular depth estimation, and NYU-Depth v2 (2012) fir surface normal estimation. For semantic segmentation, we assign labels by identifying the nearest neighbor RGB color value, and accuracy is evaluated using the Mean Intersection over Union (mIoU) metric. For monocular depth estimation, we average the output image across the three channels and apply the inverse of the linear transformation used during training to obtain depth estimates within the range of $[0,10]$ meters. Accuracy is evaluated using the Root Mean Square Error (RMSE). For surface normal estimation, we recover the corresponding normal vectors from the output image and use the Mean Angle Error to assess accuracy.
@@ 4.3 Third Section Results (Image Generation)
Settings. In this section, we focus on evaluating the effectiveness of PixWizard’s generation capabilities. Specifically, we assess its performance across four key tasks: classic text-to-image generation, controllable image generation, inpainting, and outpainting.
For controllable image generation, we evaluated PixWizard’s ability to generate images based on two specific conditions: canny edge maps and depth maps. Following ControlNet++ (Li et al., 2024), we assessed controllability by measuring the similarity between the input conditions and the corresponding features extracted from the generated images. Specifically, we used RMSE for depth map control and F1-Score for canny edges. Additionally, to evaluate the quality of the generated images and their alignment with the input text, we reported FID (Fréchet Inception Distance) and CLIP-Score metrics. All experiments were conducted at a resolution of $512\times 512$.
For image inpainting—where masked regions of an image are filled with new content—we use latent diffusion settings (Rombach et al., 2022b) to measure FID and LPIPS, which assess the quality of the generated samples when 40-50% of the image area needs to be inpainted. For image extrapolation (outpainting), we follow MaskGIT (Chang et al., 2022a) settings, extending the image by 50% to the right and comparing performance against common baselines using FID and Inception Score (IS). Both tasks are evaluated on 30,000 image crops, each sized $512\times 512$, from the Places dataset (Zhou et al., 2017a).
@@
Table 3: Comparison of PixWizard with task-specific and vision generalist baselines across six representative tasks, covering both high-level visual understanding and low-level image processing. ’$\times$’ indicates that the method is incapable of performing the task
@@
Figure 9: Visualization examples under different conditions.
Inpainting Results. The comparison with other inpainting approaches in Table 3 shows that PixWizard improves overall image quality, as indicated by FID and LPIPS. We attribute this to PixWizard’s ability to "paint" on a blank canvas, which greatly enhances its capacity to identify masked areas and generate coherent content within them. With this inpainting capability, PixWizard allows for more precise image editing tasks: (i) Remove Anything tackles the object removal problem (Criminisi et al., 2003; 2004; Elharrouss et al., 2020), enabling users to remove specific objects while keeping the result visually seamless. The process consists of two steps: identifying and removing. As shown in Fig. 10, the user first instructs PixWizard on which object to locate, and PixWizard generates a target mask. The mask is then applied to the original image, and PixWizard fills the area with appropriate background details. (ii) Replace Anything lets users swap out any object in an image. The process is similar to Remove Anything, but instead of just removing the object, the model replaces it with the specified object while ensuring the background remains visually consistent. (iii) Add Anything enables users to insert any object into an image and place it wherever they like. The user adds a mask to the desired area and provides a text prompt specifying what to add. PixWizard then leverages its powerful inpainting ability to generate the requested content.
@@ 4.4 Ablation Study
In the Sec. D.3, we conduct ablation studies to examine how the design of each component and the training process impact learning.
@@ B.2 Image Grounding
Human Pose to Image. For human pose-based generation, we employed the OpenPose model (Cao et al., 2017) for real-time multi-person 2D pose estimation. To ensure quality, only images where at least 30% of the key points of the whole body were detected were retained; those with fewer detected key points were discarded. We directly use visualized pose images with human skeletons as input condition. Finally, we obtain 0.4M pose-image-caption pairs.
@@ D.2 Image Grounding
D.3 Ablation Study
We performed ablation studies to assess the impact of each component’s design and the training process on learning in PixWizard. Given computational limitations, we conducted the ablation on PixWizard, training it for 40k steps.
Methods Rain100L val NYUv2 Emu Edit MultiGen-20M Places
PSNR↑ cIoU↑ RMSE↓ $\text{CLIP}_{dir}\!\uparrow$ F1↑ FID↓ FID↓
PixWizard w/o DSTS 30.19(-0.05) 41.66(-0.06) 0.326(-0.002) 0.091(+0.003) 33.12(+0.11) 16.98(+0.16) 10.94(-0.06)
PixWizard w/o two-stage 29.17(-1.07) 40.18(-1.54) 0.322(+0.002) 0.085(-0.003) 33.07(+0.06) 17.21(-0.07) 11.07(-0.19)
PixWizard 30.24 41.72 0.324 0.088 33.01 17.14 10.88
Influence of Dynamic Semantic Tokens Sampling. As shown in Table 6, the impact of using or not using Dynamic Semantic Tokens Sampling on task performance is minimal. This suggests that it’s unnecessary to model semantic tokens for the entire image. Since different tasks focus on different semantic areas, dynamically sampling the required features allows the model to handle each task more efficiently. Additionally, using fewer tokens reduces the computational load during attention calculations, resulting in faster inference times.

### v2

@@ Abstract
This paper presents a versatile image-to-image visual assistant, PixWizard, designed for image generation, manipulation, and translation based on free-from language instructions. To this end, we tackle a variety of vision tasks into a unified image-text-to-image generation framework and curate an Omni Pixel-to-Pixel Instruction-Tuning Dataset. By constructing detailed instruction templates in natural language, we comprehensively include a large set of diverse vision tasks such as text-to-image generation, image restoration, image grounding, dense image prediction, image editing, controllable generation, inpainting/outpainting, and more. Furthermore, we adopt Diffusion Transformers (DiT) as our foundation model and extend its capabilities with a flexible any resolution mechanism, enabling the model to dynamically process images based on the aspect ratio of the input, closely aligning with human perceptual processes. The model also incorporates structure-aware and semantic-aware guidance to facilitate effective fusion of information from the input image. Our experiments demonstrate that PixWizard not only shows impressive generative and understanding abilities for images with diverse resolutions but also exhibits promising generalization capabilities with unseen tasks and human instructions. The code and related resources are available at https://github.com/AFeng-x/PixWizard.
@@ 1 Introduction
$\bullet$ 1. Task Unification: Given the diverse nature of vision tasks and data formats, ranging from pixels and coordinates to binary masks and categories, it is challenging to find a unified representation. We observe that most vision tasks can be framed as image-to-image translation problems. For tasks not naturally suited to image output, we first learn to generate their visualizations and then apply post-processing to convert them into the desired formats. This approach is a key step toward developing a versatile visual assistant.
$\bullet$ 2. Data Construction: We aim to leverage and integrate the remarkable diversity of tasks and data in the visual domain. To achieve this, we have built a comprehensive training set with a total of 30 million data points, enabling our model to support five main capabilities: (i) Image generation, which includes text-to-image generation, controllable generation, inpainting, and outpainting; (ii) Image editing; (iii) Image restoration, covering tasks such as deraining, desnowing, deblurring, super-resolution, and more; (iv) Image grounding, which involves locating objects based on user prompts; and (v) Dense image prediction, which includes depth estimation, surface normal estimation, pose estimation, semantic segmentation, and image-to-canny/HED/sketch/Line-Art conversions.
$\bullet$ 3. Architecture Design: For a robust visual assistant, the architecture and scalability of the foundational model are crucial. In this work, we use the flow-based Diffusion Transformer (DiT) (Ma et al., 2024) as our base model. Compared to traditional diffusion models, flow-based models offer greater versatility and stability in modeling data distributions by learning conditional velocity fields. The DiT architecture further enhances scalability and is well-suited for incorporating conditional information. Building on this foundation, we introduce several unique and effective components. Specifically, we extend the dynamic partitioning and padding scheme to enable the model to handle input images of any resolution, closely aligning with human perceptual processes. Additionally, we introduce structure-aware and semantic-aware guidance, allowing the model to effectively follow multimodal instructions (images and open-ended user prompts) for a variety of manipulations.
@@ 2 Omni Pixel-to-Pixel Instruction-Tuning Dataset
To equip our image-to-image visual assistant with comprehensive capabilities in image generation, manipulation, and translation, we compiled a multi-task training dataset for visual instruction tuning, consisting of 30 million instances across seven primary domains, as illustrated in Fig. 1. To our knowledge, this is the largest and user-friendly image-instruction-image triplet dataset, built from both open-source and in-house data, filtered with the help of MLLMs and manual review.
Image Restoration. We incorporate low-level data to restore images degraded by various environmental or technical factors. This section utilizes a wide range of open-source datasets covering key restoration tasks, including (1) Denoising, (2) Deraining, (3) Demoireing, (4) Dehazing, (5) Deblurring, (6) Desnowing, (7) Deshadowing, (8) Low-Light Enhancement, (9) Face Restoration, (10) Watermark Removal, and (11) Super Resolution. Since both the inputs and outputs are inherently defined in the RGB space, these tasks can be seamlessly unified by our PixWizard model without extra any transformations. All open-source datasets we use are provided in Sec. B.1.
@@
Inpainting involves filling in missing parts of an image with new or modified content. To create inpainting instances, we apply random black or white masks to different regions of the original images. These masks come in various shapes, including circles, rectangles, and free-form patterns, and are randomly placed on the images, resulting in a wide range of occluded areas for inpainting.
@@ 3.2 Architecture
Semantic-Aware Guidance. Besides recognizing structural features, we use CLIP L/14-336 (Radford et al., 2021) to obtain semantic image embeddings. Within the PixWizard block, we introduce two zero-initialized attention mechanisms, allowing latent target image tokens to query information from the condition keys and values. Specifically, a zero-initialized gating mechanism is employed to gradually inject conditional image and text information into the token sequences. Given target image queries ($Q_{i}$), keys ($K_{i}$), and values ($V_{i}$), along with text instruction keys ($K_{t}$) and values ($V_{t}$), and conditional image keys ($K_{ci}$) and values ($V_{ci}$), the final attention output is formulated as:
@@ 3.3 Any Resolution
PixWizard inherits the dynamic partitioning and padding scheme introduced by (Zhuo et al., 2024), allowing the model to handle images of any resolution and aspect ratio during both fine-tuning and inference. However, in practice, the required resolutions for different tasks can vary significantly. To support more flexible handling of arbitrary resolutions while preserving the original resolution as much as possible, we use $[512^{2},768^{2},1024^{2}]$ as resolution centers to generate candidate patch partitions. During training, we group data with similar resolutions into buckets, ensuring that sequence lengths within each batch are comparable, minimizing padding tokens, and improving training efficiency. During inference, by incorporating NTK-Aware Scaled RoPE (Peng & Quesnelle, 2023) and sandwich normalization, PixWizard demonstrates exceptional resolution extrapolation.
@@ 4.1 First Section Results
Dense image prediction tasks are evaluated on the three vision tasks: ADE20k (2017b) for semantic segmentation, NYUv2 (2012) and SUNRGB-D (2014) for monocular depth estimation, and NYU-Depth v2 (2012) for surface normal estimation. For semantic segmentation, we assign labels by identifying the nearest neighbor RGB color value, and accuracy is evaluated using the Mean Intersection over Union (mIoU) metric. For monocular depth estimation, we average the output image across the three channels and apply the inverse of the linear transformation used during training to obtain depth estimates within the range of $[0,10]$ meters. Accuracy is evaluated using the Root Mean Square Error (RMSE). For surface normal estimation, we recover the corresponding normal vectors from the output image and use the Mean Angle Error to assess accuracy.
@@ 4.3 Third Section Results (Image Generation)
In this section, we focus on evaluating the effectiveness of PixWizard’s generation capabilities. Specifically, we assess its performance across four key tasks: classic text-to-image generation, controllable image generation, inpainting, and outpainting.
Settings. For controllable image generation, we evaluated PixWizard’s ability to generate images based on two specific conditions: canny edge maps and depth maps. Following ControlNet++ (Li et al., 2024), we assessed controllability by measuring the similarity between the input conditions and the corresponding features extracted from the generated images. Specifically, we used RMSE for depth map control and F1-Score for canny edges. Additionally, to evaluate the quality of the generated images and their alignment with the input text, we reported FID (Fréchet Inception Distance) and CLIP-Score metrics. All experiments were conducted at a resolution of $512\times 512$.
For image inpainting, we use latent diffusion settings (Rombach et al., 2022b) to measure FID and LPIPS, which assess the quality of the generated samples when 40-50% of the image area needs to be inpainted. For image extrapolation (outpainting), we follow MaskGIT (Chang et al., 2022a) settings, extending the image by 50% to the right and comparing performance against common baselines using FID and Inception Score (IS). Both tasks are evaluated on 30,000 image crops, each sized $512\times 512$, from the Places dataset (Zhou et al., 2017a).
@@
Table 3: Comparison of PixWizard with task-specific baselines across five representative tasks.
@@
Figure 9: Visualization examples under different conditions.
@@
Figure 10: Visualization results of Inpainting and Outpainting.
Inpainting Results. The comparison with other inpainting approaches in Table 3 shows that PixWizard improves overall image quality, as indicated by FID and LPIPS. The qualitative examples in Fig. 10 further demonstrate PixWizard’s effectiveness in generating coherent content. We attribute this to PixWizard’s ability to "paint" on a blank canvas, which greatly enhances its capacity to identify masked areas and generate coherent content within them.
Building on this strong inpainting capabilities,, PixWizard allows for more precise image editing tasks: (i) Remove Anything tackles the object removal problem (Criminisi et al., 2003; 2004; Elharrouss et al., 2020), enabling users to remove specific objects while keeping the result visually seamless. The process consists of two steps: identifying and removing. As shown in Fig. 11, the user first instructs PixWizard on which object to locate, and PixWizard generates a target mask. The mask is then applied to the original image, and PixWizard fills the area with appropriate background details. (ii) Replace Anything lets users swap out any object in an image. The process is similar to Remove Anything, but instead of just removing the object, the model replaces it with the specified object while ensuring the background remains visually consistent. (iii) Add Anything enables users to insert any object into an image and place it wherever they like. The user adds a mask to the desired area and provides a text prompt specifying what to add. PixWizard then leverages its powerful inpainting ability to generate the requested content.
@@ 4.4 Ablation Study
In this section, we performed ablation studies to assess the impact of each component’s design and the training process on learning in PixWizard. Given computational limitations, we conducted the ablation on PixWizard, training it for 40k steps.
Methods Rain100L val NYUv2 Emu Edit MultiGen-20M Places
PSNR↑ cIoU↑ RMSE↓ $\text{CLIP}_{dir}\!\uparrow$ F1↑ FID↓ FID↓
PixWizard w/o DSTS 30.19(-0.05) 41.66(-0.06) 0.318(+0.006) 0.091(+0.003) 32.93(-0.08) 17.02(+0.12) 10.93(-0.05)
PixWizard w/o two-stage 29.17(-1.07) 40.18(-1.54) 0.322(+0.002) 0.085(-0.003) 32.87(-0.14) 17.21(-0.07) 10.97(-0.09)
PixWizard 30.24 41.72 0.324 0.088 33.01 17.14 10.88
Influence of Dynamic Semantic Tokens Sampling. As shown in Table 4, the impact of using Dynamic Semantic Tokens Sampling (DSTS) on task performance is minimal. However, on average, using the DSTS method leads to better overall performance. This suggests that it’s unnecessary to model semantic tokens for the entire image. Since different tasks focus on different semantic regions, dynamically sampling the relevant features allows the model to handle each task more efficiently. Additionally, using fewer tokens reduces the computational load during attention calculations, leading to faster inference times.
@@ B.2 Image Grounding
Human Pose to Image. For human pose-based generation, we employed the OpenPose model (Cao et al., 2017) for real-time multi-person 2D pose estimation. To ensure quality, only images where at least 30% of the key points of the whole body were detected were retained; those with fewer detected key points were discarded. We directly use visualized pose images with human skeletons as input condition. Finally, we obtain 0.25M pose-image-caption pairs.